IT Software Project

As a senior analyst for the company you have been asked to evaluate a new IT software project. The company has just paid a consulting firm $50,000 for a test marketing analysis. After looking at the project plan, you anticipate that the project will need to acquire computer hardware for a cost of $400,000. The Australian Taxation Office rules allow an effective life for the computer hardware of five years. The equipment can be depreciated on a straight-line (prime cost) basis and there is no expected salvage value after the five years.

Your company does not have any available space where the project can be located for five years and you anticipate a new office will cost $80,000 to rent for the first year. You expect that the project will need to hire 3 new software specialists at $50,000 (each specialist) in the first year for the full five years to work on the software.

The project will use a van currently owned by the company and although the van is not currently being used by the company, it can be rented out for $5,000 per year for five years (inclusive inflation). The book value of the van is $20,000. The van is being depreciated straight-line (with five years remaining for depreciation) and is expected to be worthless after the five years.

Expected annual marketing and selling costs will be incurred during the life of the project (5 years), with the first year expecting to be $200,000. The produced software is expected to sell at $100 per unit while the cost to produce each unit is $40. You expect that 10,000 units will be sold in the first year and the number of units sold will increase by 20% a year for the remaining four years. The project will need working capital of $50 000 to commence the business (in year 0) and the investment in working capital is to be completely recovered by the end of the project’s life (in year 5). The company tax rate is 30%, and the discount rate is 10%.

Based on the information presented above, answer the following questions (1) – (3).

  1. (8 marks)

In evaluating the new IT software project, are the cost of $50,000 spent on marketing analysis and the use of van relevant for capital budgeting decision? Explain your answer(s).

  1. (12 marks)

Calculate the incremental free cash flow during the project’s life (at the end of Years 1 through 5). Show workings.

  1. (10 marks)

Calculate the NPV, payback period and IRR of the project. Should the project be accepted? Show workings and explain your answer(s).
Answer #1

question No 1.

In evaluating the new IT software project, are the cost of $50,000 spent on marketing analysis and the use of van relevant for capital budgeting decision? Explain your answer(s).

  1. Cost total $50,000 spent on marketing analysis is relevant in capital budgeting since it will come under Marketing and advertising / research expense which is also required for budgeting.
  2. The use of van is relevant since company is going to use the van which is already rented so it will definitely impact capital budget.

question No 2

Calculate the incremental free cash flow during the project’s life (at the end of Years 1 through 5). Show workings.

Answer: The incremental cash flow =incremental sales - incremental operating expenses + changes in noncash operating expenses (Depreciation)

Incremental Sales:

Year 1 10000*100$ = 1000000

Year 2 12000*100$ = 1200000

Year 3 14400*100$ = 1440000

Year 4 17280*100$=1728000

Year 5 20736*100$=2073600

Operation Expense is same for years. So 40per unit

Year 1 10000*40$ = 400000

Year 2 12000*40$ = 480000

Year 3 14400*40$ = 576000

Year 4 17280*40$=691200

Year 5 20736*40$=829440

Depreciation expense for van and computer

Year 84000$ For van and Computer

So Answer is

Year 1 1000000-400000+84000 = 684000

Year 2 1200000-480000+84000=804000

Year 3 1440000-576000+84000= 948000

Year 4 1728000-691200+84000=1120800

Year 5 20736000-829440+84000=1328160

Question no 3.

Calculate the NPV, payback period and IRR of the project. Should the project be accepted? Show workings and explain your answer(s).

NPV= NPV = (Cash flows)/( 1+r)i

Cash flows= Cash flows in the time period

r = Discount rate

i = time period

intial Investment = 400000 Computer +20000 van +50000 intial cash

470000

Year

Flow

Present value

Computation

0

-470000

-470000

1

684000

621818.18

684000/(1.1)

2

804000

664462.8

804000/(1.1)^2

3

948000

712246.43

948000/(1.1)^3

4

1120800

765521.4

1120800/(1.1)^4

5

1328160

824682.8

1328160/(1.1)^5

Pay Back period is 1 Year

(Already answered 4 Subparts)

IRR is ~160.58% calculated online with initial investment as 470000$

The project should be accepted.